# Merging Facts, Crafting Fallacies: Evaluating the Contradictory Nature of Aggregated Factual Claims in Long-Form Generations

### 08 Feb 2024 v1

@@ 4.1 Definition
D-FActScore is based on two key concepts to handle entity ambiguity properly. The following two key concepts mark the most significant difference between D-FActScore and prior factuality metrics (Min et al., 2023a; Gao et al., 2023).
Key concept 1: For a factual paragraph, atomic facts about distinct entities must be distinguishable. A long-form generation may include the facts of different entities. If two atomic facts from the same paragraph refer to different entities, a factual paragraph must allow the readers without prior knowledge to understand that the two facts belong to distinct entities. If the paragraph does not provide sufficient information for readers to see that the two facts are about different entities, the readers can falsely believe that there exists a single entity that satisfies the two facts.
Key concept 2: The atomic facts for the same individual should be non-conflicting. The concept of same individual is related to key concept 1. Readers will consider two mentions to be the same individual if the paragraph does not provide enough information to convince readers that the two mentions correspond to different entities. FActScore and citation recall consider facts, and two facts about the same individual may be verified using the information of different entities. However, atomic facts referring to the same individual are interrelated and thus should be considered together.
Instead of considering the factuality of each atomic fact independently based on a given knowledge source, D-FActScore first groups the atomic facts referring to the same individual into a fact group. Next, for each fact group, D-FActScore finds an entity in the knowledge source that best matches the facts in that fact group. We call the entity found by entity linking the "linked entity". Last, for all the facts in the same fact group, we verify whether they can be supported based on the linked entity’s information in the knowledge source.
Definition. Let $y$ be a generated paragraph from an LLM. Let $\mathcal{A}_{y}$ be a list of atomic facts decomposed from $y$. Based on the above key concepts, we group the atomic facts into $N$ groups of atomic facts $\{\mathcal{A}_{y,1},\cdots,\mathcal{A}_{y,N}\}$. Two atomic facts belong to the same group if they are about the same individual. For each group of atomic fact $\mathcal{A}_{y,i}$, we use entity linking to find its corresponding entity $e_{i}^{*}$ in the knowledge source $\mathcal{C}$. Let the subset of the knowledge source related to $e_{i}^{*}$ denoted by $C_{i}^{*}$. The D-FActScore of $y$ is defined as follows:
@@
In our paper, we use the paragraphs generated in Section 2 as $y$ and use Wikipedia as $\mathcal{C}$.
@@ 4.2 Annotating D-FActScore by Humans
We conduct human evaluation to calculate the D-FActScore of paragraphs generated in Section 2. The annotators are presented with a paragraph, the atomic facts decomposed from the paragraph, and all the Wikipedia pages of the entities with the target name. The annotations are conducted via the following steps.
@@
Step 2: Determine the number of bios. We instruct the annotators to determine the number of biographies based on the paragraph, neglecting any prior knowledge or the Wikipedia pages we prepare. Identifying more than one biography indicates that the paragraph provides enough information to separate the biographies of distinct entities. We also ask the annotators to link each biography to a Wikipedia page we present.
Step 3: Label factual precision. For each atomic fact, check from which biography the atomic fact is extracted; that biography is linked to a Wikipedia page in Step 2. Next, determine if the fact is Irrelevant to the target name. If it is not Irrelevant, verify whether the atomic fact is Supported or Not-Supported by the Wikipedia page linked to the biography.
Determining the number of distinguishable biographies is somewhat subjective since some people consider a narrative to provide enough disambiguation information while some do not. Despite this subjectivity, the annotators agree highly on the number of biographies in the paragraph.
Three annotators from Upwork are hired to label 300 paragraphs generated in Section 2 from Llama-13b-chat, Tulu-v2-13b-dpo, and ChatGPT; 100 paragraphs per model. We choose these models since they include open-source and proprietary models and have different alignment training methods. The exact instructions, annotation interface, and agreement rate between annotators are shown in Appendix B.
@@ 4.3 Human Evaluation Results
Aside from the annotation of D-FActScore elaborated above, we ask the annotators to perform additional annotation to allow us to calculate the number of distinct entities and the FActScore of the paragraph. FActScore is annotated based on the definition in Min et al. (2023a). The number of distinct entities is calculated by performing entity linking to each atomic fact independently without considering the context and counting the number of unique Wikipedia pages linked. Details are in Appendix B.1.
D-FActScore’s evaluation is closer to how humans consider the factuality of a long-form generation and thus can be considered as the paragraph’s ground truth factuality. The human annotation results of D-FActScore and FActScore are presented in Table 2(a). We have the following observations.
This is because all models mix the facts of multiple entities in a single biography, and FActScore considers these misleading paragraphs factual as long as atomic facts can be supported by Wikipedia. D-FActScore does not have this problem by construction. The gap between FActScore and D-FActScore can be interpreted as the tendency of an LLM to mix the facts of multiple entities in an ambiguous way.
FActScore and D-FActScore yield different model rankings.
@@
Table 3: FS: FActScore, D-FS: D-FActScore, # bio: average number of separable biographies in one paragraph, # ent.: average number of distinct entities in one paragraph, chat-13b: Llama-13b-chat. Human and automatic evaluations are conducted on the same set of paragraphs.
@@
ChatGPT has the highest D-FActScore compared to the other two models, and the average number of biographies and entities is almost the same. This implies that ChatGPT can distinguish entities with the same name in the retrieved passages and generate a paragraph that provides the readers with sufficient information to disambiguate those entities.
@@ 5 Automatic Evaluation of D-FActScore
Human evaluation is time-consuming and expensive. Hence, we devise an automatic pipeline to estimate the D-FActScore (Section 5.1) and show that it can approximate the D-FActScore obtained by human annotation (Section 5.2). We then use the automatic pipeline to evaluate the generation from five LLMs (Section 5.3).
@@ 5.1 Automatic Evaluation
Step 3: Label factual precision. We need to verify all the facts in one biography based on the Wikipedia page of a single entity. We perform entity linking to find a Wikipedia entity that best matches the entity of the biography, called the linked entity. Details of entity linking are elaborated in Appendix C.1. For each biography and the facts in it, we verify those facts based on the Wikipedia page of the linked entity. Precisely, we use Retrieve $\to$ LM in Min et al. (2023a) and prompt ChatGPT to determine if an atomic fact is supported by the Wikipedia page of that entity.
@@ 5.2 Human VS Automatic Evaluation
We compare the result of D-FActScore obtained with human and automatic evaluation in Table 3.
Automatic and human evaluation of D-FActScore shows the same model ranking.   We find the factuality ranking based on D-FActScore among the three models in Table 2(b) agrees with the ranking in Table 2(a): ChatGPT is the most factual, and Llama-13b-chat is the least factual. This shows that the automatic evaluation of D-FActScore provides a reliable estimation of the relative factuality of different LLMs.
D-FActScore obtained with automatic evaluation is higher than human evaluation.   Compared with the human evaluation result of Table 2(a), D-FActScore obtained using automatic evaluation is higher, and the absolute error is at most 8%. This observation also aligns with Min et al. (2023a), which shows that using ChatGPT+retrieval to estimate FActScore can yield a higher FActScore compared to human evaluation.
Automatic evaluation can determine the number of biographies accurately.   The correctness of D-FActScore’s automatic evaluation strongly depends on the second step: splitting atomic facts into groups corresponding to biographies of different individuals. The number of groups is the number of distinguishable biographies. By comparing the number of bios obtained using automatic and human evaluation in Table 3, we find that GPT3.5 can accurately determine how many biographies there are in a paragraph. The difference between the number of biographies obtained by human evaluation and automatic evaluation differs within 0.1 in all three models, justifying using GPT3.5 to split the paragraph into biographies.
@@
5.3 Comparing Different LLMs and Different Demonstrations
After showing that D-FActScore can be estimated using automatic evaluation, we use automatic evaluation to evaluate the D-FActScore of paragraphs generated using all 500 names in AmbigBio by five LLMs and two types of demonstrations. We have the following findings.
Demonstrations with name ambiguity do not make the outputs more factual.   In Table 4, we compare the results obtained with two types of demonstrations: names with and without ambiguity. Recall that in the demonstration with name ambiguity, the target name is a name with ambiguity, and the top-5 retrieved documents contain passages from different entities with the same name. Since the demonstration answer only contains a biography that includes the information of one entity, we hope the LLM can know how to handle target names with ambiguity better. However, we do not see a higher D-FActScore when using demonstrations with name ambiguity.
Open-source LLMs lag behind ChatGPT.   All the open-sourced models we use have a D-FActScore much lower than that of ChatGPT. Furthermore, the entity per paragraph is higher than biography per paragraph for open-sourced models, showing that the open-sourced models cannot distinguish entities with name ambiguity and generate a factual paragraph. This highlights a potential direction of improvement for open-source LLMs.
@@ 5.3 Comparing Different LLMs and Different Demonstrations
ChatGPT fully uses the retrieved documents.   By examining the passages retrieved using GTR, we estimate that the top-5 passages provided to the LLM contain 2.2 distinct entities with the target name on average. Meanwhile, the passages generated by ChatGPT also contain around 2.2 biographies and 2.2 distinct entities on average. This indicates that ChatGPT can include diverse information that is presented in the retrieved passages. This is a desirable behavior since the answer to an ambiguous question should cover the answer to as many disambiguated questions Min et al. (2020); Stelmakh et al. (2022).
The central figure of the biography is not always the most common Wikipedia entity.   When multiple entities with the target name are included in the top-5 retrieved passages, the LLM can generate a biography for any of them. We want to answer the question: When there is only one biography in the generated paragraph, but there are multiple valid entities in the retrieved passages, is the central figure of the biography always the most common entity among the retrieved entities? We hypothesize that popular entities are more likely to be included in the LLM’s training data multiple times, making the LLM more familiar with those common entities and prone to use them as the central figure. We assess how common an entity is based on the page view of their Wikipedia within the past year. However, for all LLMs, we find that the LLM picks the most popular entity to write a biography for in only 45% to 55% of the cases. This shows that an entity’s popularity does not strongly affect which entity the LLM picks to generate a biography for.
@@ 7 Conclusion and Discussion
This paper focuses on an interesting phenomenon when combining factual information yields a non-factual paragraph. This fallacy can happen due to entity ambiguity. We first show that such non-factual paragraphs can be easily generated by LLMs when the retrieved contents contain multiple entities with the same name. Next, we reveal that current evaluation metrics on fine-grained factual precision cannot properly assess the factuality of these non-factual paragraphs. We modify FActScore into D-FActScore by focusing on entity disambiguation. We conduct human evaluation of D-FActScore to compare the factuality of different LLMs, and we further show that our proposed pipeline for automatic evaluation of D-FActScore aligns with human evaluation results to a reasonable extent. Using D-FActScore, we show that four open-sourced LLMs cannot properly handle entity ambiguity in the retrieved contexts to form a factual generation. We encourage future researchers to use AmbigBio and evaluate the factuality of generated paragraphs using D-FActScore.
@@
In our study, we focus on evaluating D-FActScore using human biographies. We justify the reasons for doing this in Section 2. In fact, the difficulty of factuality evaluation due to entity ambiguity can happen in more diverse contents, and the core concepts and evaluation procedure of D-FActScore are general and can be applied to more diverse contents.
The knowledge source used in retrieval
Our findings highlight LLMs’ challenges in producing factual content with entity ambiguity when the retrieved contents come from Wikipedia. This is a simplified scenario since Wikipedia is not the only source to provide information about entities. Notably, even when provided with ample disambiguating information, i.e., the title of the Wikipedia pages of different entities, open-source models struggle to differentiate between entities accurately. This issue will become more pronounced in realistic scenarios where LLMs access content from sources beyond Wikipedia, increasing the difficulty of correctly identifying distinct entities. Addressing this challenge of entity ambiguity is crucial for deploying trustworthy retrieval-augmented LLMs.
@@
Despite the limitations listed above, we believe that our paper significantly contributes to identifying a realistic evaluation problem that has never been studied and proposes a simple solution.
@@ Appendix A Generating Paragraphs from LLMs
For all open-sourced models, we use a temperature of $1$ and top-p sampling with $p=0.95$.

### 23 Feb 2024 v2

@@ 4.1 Definition
We first define D-FActScore and then present the key concepts that motivate its definition.
Definition.  Let $y$ be a generated paragraph from an LLM. Let $\mathcal{A}_{y}$ be a list of atomic facts decomposed from $y$. We split $\mathcal{A}_{y}$ into $N$ disjoint fact groups $\{\mathcal{A}_{y,1},\cdots,\mathcal{A}_{y,N}\}$. For two atomic facts $a,a^{\prime}\in\mathcal{A}_{y}$, they will be grouped into the same fact group $\mathcal{A}_{y,i}$ if a reader without prior or external knowledge may think that $a$ and $a^{\prime}$ are about the same individual when reading the paragraph. We use the term individual to refer to a character perceived by the reader; this is different from an entity that exists in the real world. For example, all the atomic facts in Figure 1 are grouped into a single fact group since the paragraph looks like it is about the same individual named Dick Hanley.
For each group of atomic fact $\mathcal{A}_{y,i}$, we use entity linking to find an entity $e_{i}^{*}$ in the knowledge source $\mathcal{C}$ that best matches the facts in $\mathcal{A}_{y,i}$. Let the subset of the knowledge source related to $e_{i}^{*}$ denoted by $C_{i}^{*}$. For all atomic facts $a\in\mathcal{A}_{y,i}$, they will be verified using $C_{i}^{*}$ instead of using the whole $C$. The D-FActScore of $y$ is defined as follows:
@@
In our paper, $y$ is a paragraph generated in Section 2, $\mathcal{C}$ is Wikipedia, and $C_{i}^{*}$ is the Wikipedia page of $e_{i}^{*}$.
D-FActScore differs from FActScore in two aspects. Difference 1: D-FActScore groups atomic facts by their originating paragraph before verifying them. Difference 2: D-FActScore restricts that all the facts in the same group must be verified using the information of the same entity.
Motivation.  Atomic facts from a paragraph often have connections, so evaluating their factuality should consider these relationships, rather than verifying each fact independently as FActScore does. Grouping atomic facts helps manage these connections: Facts within the same group look like they are about the same individual to the readers, so their truthfulness should be confirmed using the same entity in the knowledge source. For example, the atomic facts in Figure 1 belong to one group, and they can only be supported by either the Wikipedia page of Dick Hanley (Swimmer) or Dick Hanley (American Football), but not both.
By organizing facts into groups and limiting the source of fact verification for each group, D-FActScore assesses the factuality of a paragraph with entity ambiguity more accurately. These two key differences mark the most significant difference between D-FActScore and FActScore, enabling D-FActScore to handle entity ambiguity.
@@ 4.2 Annotating D-FActScore by Humans
We conduct human evaluation to calculate the D-FActScore of paragraphs generated in Section 2. The annotators are presented with a paragraph, the atomic facts decomposed from the paragraph, and all the Wikipedia pages of the entities with the target name. The exact instructions, annotation interface, and agreement rate between annotators are shown in Appendix B. The annotations are conducted via the following steps.
@@
Step 2: Determine the number of bios. We instruct the annotators to determine the number of biographies based on the paragraph, neglecting any prior knowledge or the Wikipedia pages we prepare. Identifying more than one biography indicates that the paragraph provides enough information to separate the biographies of distinct individuals. This step essentially splits atomic facts into groups since atomic facts of the same biography are about the same individual, so they fall into the same group. We also ask the annotators to link each biography to an entity’s Wikipedia page we present. Even though determining the number of biographies in a paragraph is based on personal interpretation, we find that annotators reach a high level of agreement on the number of bios.
Step 3: Verifying atomic facts. For each atomic fact, the annotators first check if the fact is Irrelevant to the target name. If it is not Irrelevant, verify whether the atomic fact is Supported or Not-Supported using one entity’s Wikipedia page. Recall that the atomic facts in the same group belong to one of the biographies in the original paragraph and the biography is linked to an entity’s Wikipedia page in Step 2. All the atomic facts in the same group are verified using the same linked entity’s Wikipedia page.
We calculate the D-FActScore based on the human annotation results. Three annotators from Upwork are hired to label 300 paragraphs generated in Section 2 from Llama-13b-chat, Tulu-v2-13b-dpo, and ChatGPT; 100 paragraphs per model. We choose these models since they include open-source and proprietary models and have different alignment training methods.
@@ 4.3 Human Evaluation Results
Aside from the annotation of D-FActScore elaborated above, we ask the annotators to perform additional annotation to allow us to calculate the number of distinct entities and the FActScore of the paragraph. FActScore is annotated based on the definition in Min et al. (2023a). The number of distinct entities is calculated based on the definition in Section 2.4. Details are in Appendix B.1. The human annotation results of D-FActScore and FActScore are presented in Table 2(a). We have the following observations.
@@
Table 3: Human and automatic evaluations are conducted on the same set of paragraphs. FS: FActScore, D-FS: D-FActScore, # bio: average number of separable biographies in one paragraph, # ent.: average number of distinct entities in one paragraph, chat-13b: Llama-13b-chat.
This is because all models mix the facts of multiple entities in a single biography, and FActScore considers these misleading paragraphs factual as long as each atomic fact can be supported by Wikipedia. D-FActScore does not have this problem by construction. The gap between FActScore and D-FActScore can be interpreted as the tendency of an LLM to mix the facts of multiple entities in a non-factual way.
FActScore and D-FActScore yield different model rankings.
@@
ChatGPT has the highest D-FActScore, and the average number of biographies and entities is almost the same. This implies that ChatGPT can distinguish entities with the same name in the retrieved passages and generate a factual paragraph that provides the readers with sufficient information to disambiguate those entities.
@@ 5 Automatic Evaluation of D-FActScore
Human evaluation is time-consuming and expensive. Hence, we devise an automatic pipeline to estimate D-FActScore (Section 5.1) and show that it can approximate the D-FActScore obtained by human annotation (Section 5.2). We then use the automatic pipeline to evaluate the generation from five LLMs (Section 5.3).
@@ 5.1 Automatic Evaluation
Step 3: Verifying atomic facts. We verify the factuality of facts in the same fact group using the Wikipedia page of the same entity; a fact group corresponds to one biography in the original paragraph. For each biography and its corresponding fact group, we perform entity linking to find a Wikipedia entity that best matches the individual of the biography and verify the facts in that bio based on the Wikipedia page of that entity. Precisely, we use Retrieve $\to$ LM in Min et al. (2023a) to prompt ChatGPT to verify the atomic fact based on the Wikipedia of that entity. Details of entity linking and the prompts used in automatic evaluation are elaborated in Appendix C.1.
@@ 5.2 Human VS Automatic Evaluation
We compare the result of D-FActScore obtained with human and automatic evaluation in Table 3. We have the following findings.
Automatic and human evaluation of D-FActScore shows the same model ranking.   We find the factuality ranking among the three models based on D-FActScore in Table 2(b) agrees with the ranking in Table 2(a): ChatGPT is the most factual, and Llama-13b-chat is the least factual. This shows that the automatic evaluation of D-FActScore provides a reliable estimation of the relative factuality of different LLMs.
D-FActScore obtained with automatic evaluation is higher than human evaluation.   Compared with the human evaluation result of Table 2(a), D-FActScore obtained using automatic evaluation is higher, and the absolute error is at most 8%. This observation also aligns with Min et al. (2023a), which shows that using Retrieve $\to$ LM to estimate FActScore can yield a higher FActScore compared to human evaluation.
Automatic evaluation can determine the number of biographies accurately.   The correctness of D-FActScore’s automatic evaluation strongly depends on the second step: splitting atomic facts into groups corresponding to biographies of different individuals. The number of groups is the number of distinguishable biographies. By comparing the number of bios obtained using automatic and human evaluation in Table 3, we find that GPT3.5 can accurately determine how many biographies there are in a paragraph. The difference between the number of biographies obtained by human evaluation and automatic evaluation differs within 0.1 in all three models, justifying using GPT3.5 to split the atomic facts into fact groups.
5.3 Different LLMs and Demonstrations
After showing that D-FActScore can be estimated using automatic evaluation, we use automatic evaluation to evaluate the D-FActScores of paragraphs generated using all 500 names in AmbigBio by five LLMs and two types of demonstrations. We have the following findings.
Demonstrations with name ambiguity do not make the outputs more factual.   In Table 4, we compare the results obtained with two types of demonstrations: names with and without ambiguity. Recall that in the demonstration with name ambiguity, the target name is a name with ambiguity, and the top-5 retrieved documents contain passages from different entities with the same name. Since the demonstration <answer> only contains a biography that includes the information of one entity, we hope the LLM can know how to handle target names with ambiguity better. However, we do not see a higher D-FActScore when using demonstrations with name ambiguity.
@@ 5.3 Different LLMs and Demonstrations
Open-source LLMs lag behind ChatGPT.   All the open-source models we use have a D-FActScore much lower than that of ChatGPT. Furthermore, the entity per paragraph is higher than biography per paragraph for open-source models, showing that the open-source models cannot distinguish entities with name ambiguity and generate a factual paragraph. This highlights a potential direction of improvement for open-source LLMs.
@@
ChatGPT fully uses the retrieved documents.   By examining the passages retrieved by GTR, we estimate that the top-5 passages provided to the LLM contain 2.2 distinct entities with the target name on average. Meanwhile, the passages generated by ChatGPT also contain around 2.1 to 2.3 biographies and distinct entities on average. This indicates that ChatGPT can include diverse information presented in the retrieved passages. This is a desirable behavior since the answer to an ambiguous question should cover the answer to as many disambiguated questions Min et al. (2020); Stelmakh et al. (2022).
The central figure of the biography is not always the most common Wikipedia entity.   When multiple entities with the target name are included in the top-5 retrieved passages, the LLM can generate a biography for any of them. We want to answer the question: When there is only one biography and one entity in the generated paragraph, but multiple valid entities exist in the retrieved passages, is the central figure of the biography always the most common entity among the retrieved entities? We hypothesize that popular entities are more likely to be included in the LLM’s training data more times, making LLM more familiar with those common entities and prone to use them as the central figure. We assess how common an entity is based on the page view of their Wikipedia over the past year. However, for all LLMs, we find that in only 45% to 55% of the cases, the LLM picks the most popular entity to write a biography for. This shows that an entity’s popularity does not strongly affect which entity the LLM picks to generate a biography for.
@@ 6 Related Work
We discuss the related works in the appendix due to the limited space. If the paper is accepted, we plan to move this section to the main content using the additional page in the camera-ready version.
@@ 7 Conclusion and Discussion
We show that combining factual information yields a non-factual paragraph due to entity ambiguity, and this kind of paragraph can be easily generated by LLMs using RAG. We further reveal that current factuality metrics cannot correctly assess the factuality of these non-factual paragraphs. To resolve the evaluation problem, we modify FActScore into D-FActScore by focusing on entity disambiguation. We conduct human evaluations of D-FActScore to compare the factuality of different LLMs and show that D-FActScore of open-source LLMs lag behind ChatGPT, indicating that open-source LLMs cannot handle ambiguous entities in the retrieved documents to form a factual narrative. We propose a pipeline for automatic evaluation of D-FActScore and show it aligns with human evaluation results to a reasonable extent. We encourage future researchers to use AmbigBio and evaluate the factuality of generated paragraphs using D-FActScore. The ambiguous name collection AmbigBio and the codes for D-FActScore are available at https://github.com/d223302/Merging-Facts-Crafting-Fallacies.
Our findings underscore LLMs’ difficulties in generating accurate content when retrieving from Wikipedia with entity ambiguity. The scenario in our paper is simplified, given that Wikipedia is not the sole information source about entities. Even using high-quality retrieved content from Wikipedia, open-source models struggle to differentiate between entities accurately. In more realistic situations where LLMs draw from a broader and more nuanced content pool on the Web, and the named entities used to query LLMs may not appear in the training data, distinguishing between different entities becomes even more challenging. Overcoming this issue is vital for the reliability of retrieval-augmented LLMs.
@@
In our study, we focus on evaluating D-FActScore of human biographies. We justify the reasons for doing this in Section 2. In fact, the difficulty of factuality evaluation due to entity ambiguity can happen in more diverse contents, and the core concepts and evaluation procedure of D-FActScore are general and can be applied to more diverse contents.
@@
Using GPT3.5 to extract atomic facts
In our paper, we rely on using GPT3.5 (text-davinci-003) to extract atomic facts and split atomic facts into groups (Section 5). The reason for using GPT3.5 is to match the experiment setup of Min et al. (2023a), which also relies on GPT3.5 to extract the atomic facts. Using the same model to extract atomic facts makes it easier for us to compare with them. However, GPT3.5 was recently deprecated by OpenAI, making it impossible to reproduce the results of Min et al. (2023a) and the experiment in our paper. To alleviate this issue, we show in Table 4(c) in Appendix C.2 that using ChatGPT (gpt-3.5-turbo) to extract the atomic facts and split the atomic facts into groups yields almost the same result as using GPT3.5. All our observations in Section 5 hold when using ChatGPT to extract and group the atomic facts.
Despite the limitations listed above, we believe that our paper significantly contributes to identifying a realistic evaluation problem that has never been studied and proposing a simple solution.
We do not see specific risks or harm of our paper.
@@ Appendix A Generating Paragraphs from LLMs
For all open-source models, we use a temperature of $1$ and top-p sampling with $p=0.95$.